I have 5 word problems that I need to know how to write an algebraic expression for: Hope you can help...this is for the 6th grade...
1. Tim wants to buy an AC. The AC costs $45. It costs $2 a day for electricity to run the AC. Write an expression for D days to run the AC. If Time spent $85 to cool with the AC, howm any days did he run it?
2. Tammy wants to rent a movie,membership is $15 and it costs $4 per movie. Write an expression to find the cost to rent the V videos. If Tammy spent $103 at the video store, how many videos did she rent?
3. Sam wants to buy a cell phone. He has $45 now and can save $6 a month. Write an expression to show how much savings he will have after M months. If Sam wants to buy a phone for $63, how many months will it take him to save for it?
4. Sarah bought a membership for some training sessions. The membership costs $12 a month + $10 for each training session. Write an expression to show the cost for S number of sessions. If Sarah spent $62 in March for training sessions, how many did she buy?
5. Betty wants to mail a package. The flat rate is $4 per package + $2 per oz. Write an expression to determine the cost for the package that weighs U oz. If Betty spent $88 to mail her package, how many oz did the package weigh?
I hope this isn't too much. I can usually do these kind of problems but having had a major headache all day today has made it hard for me to even think :)
Thanks for your help!

Answer by ankor@dixie-net.com(22740)

1. Tim wants to buy an AC. The AC costs $45. It costs $2 a day for electricity to run the AC.
Write an expression for D days to run the AC.
2d + 45
If Time spent $85 to cool with the AC, how many days did he run it?
2d + 45 = 85
2d = 85-45
2d = 40
d = 20 days
:
2. Tammy wants to rent a movie,membership is $15 and it costs $4 per movie. Write an expression to find the cost to rent the V videos.
4V + 15
If Tammy spent $103 at the video store, how many videos did she rent?
4V + 15 = 103
solve for V the same way we did above
:
3. Sam wants to buy a cell phone. He has $45 now and can save $6 a month.
Write an expression to show how much savings he will have after M months.
6M + 45
If Sam wants to buy a phone for $63, how many months will it take him to save for it?
6M + 45 = 63
:
4. Sarah bought a membership for some training sessions. The membership costs $12 a month + $10 for each training session.
Write an expression to show the cost for S number of sessions.
10S + 12
If Sarah spent $62 in March for training sessions, how many did she buy?
10s + 12 = 62
;
5. Betty wants to mail a package. The flat rate is $4 per package + $2 per oz.
Write an expression to determine the cost for the package that weighs U oz.
2U + 4
If Betty spent $88 to mail her package, how many oz did the package weigh?
2U + 4 = 88
;
:
You have to admit these are not that hard. Hope your head-ache is better today.